# Positional Prompt Tuning for Efficient 3D Representation Learning

### v1

@@ Introduction
With point cloud as the only modality of input, it is difficult for 3D parameter-efficient fine-tuning tasks to get prompts that are rich in linguistic significance, like a well-designed prefix prompt in natural language, which can remarkably boost the performance of generative tasks in text modality (Lester, Al-Rfou, and Constant 2021). VPT (Jia et al. 2022b) introduce a set of trainable embeddings into the input of the Transformer to solve this problem in the vision domain, and both DAPT (Zhou et al. 2024) and IDPT (Zha et al. 2023) adopt a similar manner. Simultaneously considering Adapter and Prompt, they both generate extra prompt tokens from an adapter-structure network and add these prompt tokens into Transformer input. By doing so they acquire dynamic prompts rather than static ones in VPT. As the static prompts always need a process of manual selection, the generated dynamic prompts are a compromise of this process. Nevertheless, these kinds of prompts are actually an abstraction of the existing feature embedding. Thus the input of Transformer layers with extra prompt inserted and the Transformer output can be formulated as:
@@
Where means class token, represents the extra prompts abstracted from patch embedding . attn and F represent the FFN and attention layer inside the Transformer blocks. The and represent trainable and frozen parameters, respectively. Thus the information of prompts is highly homogeneous, which is an aggregation of existing tokens, rather than the newly added language prompts in the NLP domain. We then think about whether there are some other forms of prompts that can both be dynamically adjusted and in the meantime be of physical significance And it comes to our mind that the patch tokens themselves, which are down-sampled and clustered, can be used as extra tokens for prompting. Different from prompts that are yielded from patch token embedding, we unfreeze a part of the patch encoder to introduce variety into patch token embedding. In that case, with this form of prompts, the input of Transformer layers can be formulated as:
@@
| DGCNN (Wang et al. 2019) | TOG 19 | 31.6 2.8 | 40.8 4.6 | 19.9 2.1 | 16.9 1.5 |
| OcCo (Wang et al. 2021) | ICCV 21 | 90.6 2.8 | 92.5 1.9 | 82.9 1.3 | 86.5 2.2 |
@@
| OcCo (Wang et al. 2021) | ICCV 21 | 94.03.6 | 95.92.3 | 89.45.1 | 92.44.6 |
| Point-BERT (Yu et al. 2022) | CVPR 22 | 94.6 3.1 | 96.3 2.7 | 91.0 5.4 | 92.7 5.1 |
| Point-MAE (Pang et al. 2022) | ECCV 22 | 96.3 2.5 | 97.8 1.8 | 92.6 4.1 | 95.0 3.0 |
| Point-M2AE (Zhang et al. 2022) | NeurIPS 22 | 96.8 1.8 | 98.3 1.4 | 92.3 4.5 | 95.0 3.0 |
| ACT (Dong et al. 2023) | ICLR 23 | 96.8 2.3 | 98.0 1.4 | 93.3 4.0 | 95.6 2.8 |
| VPP (Qi et al. 2023b) | NeurIPS 23 | 96.9 1.9 | 98.3 1.5 | 93.0 4.0 | 95.4 3.1 |
| I2P-MAE (Zhang et al. 2023c) | CVPR 23 | 97.0 1.8 | 98.3 1.3 | 92.6 5.0 | 95.5 3.0 |
| ReCon (Qi et al. 2023a)(baseline) | ICML 23 | 97.3 1.9 | 98.9 1.2 | 93.3 3.9 | 95.8 3.0 |
| with Parameter-Efficient Supervised Finetuning | |||||
| ReCon w/ IDPT∗ (Zha et al. 2023) | ICCV 23 | 96.92.4 | 98.30.7 | 92.84.0 | 95.53.2 |
| ReCon w/ DAPT∗ (Zhou et al. 2024) | CVPR 24 | 95.62.8 | 97.71.6 | 91.94.1 | 94.63.5 |
| ReCon w/ PPT (Ours) | - | 97.0 2.7 | 98.71.6 | 92.25.0 | 95.62.9 |
represents the added trainable adapter for more thorough feature aggregation and adjustment. In this way, both the extra sampled patch token prompts and the original patch token embedding are altered for information integration. On this manner of prompts and adapters, combined with the aforementioned positional encoding, we carried out extensive experiments, and our method achieved excellent results in all of these experiments, which will be discussed in subsequent chapters.
Pipeline of PPT
@@
The pipeline structure of our PPT is shown in Figure 3(a). The patch encoder of the pre-trained model is already empowered to abstract a high concentration of semantic information from the input point cloud, so we unfreeze part of the patch encoder to provide varieties for patch tokens and prompt tokens. Briefly, we resample the input point clouds as prompts and concatenate them with the input patch tokens before sending them into the Transformer, and trainable positional embedding layers are concatenated together as the positional information for both the prompts and the patch tokens. Before inputting into each Transformer layer, a patch adapter layer and a prompt adapter layer are adopted as a dynamic adjustment module. There are also adapter layers inside every Transformer block, specifically between the attention layer and the FFN. As the pre-trained parts of the Transformer layers are untrainable during the tuning process, an adapter inside the block effectively integrates information and dynamically adjusts it to improve performance.
Experiments
3D Real-World Object Recognition
@@
ScanObjectNN (Uy et al. 2019) is one of the most challenging 3D datasets, which covers 15K real-world objects from 15 categories. For a fair comparison, we report the results with and without the voting strategy (Liu et al. 2019) separately. The results of object classification tasks of ScanObectNN (OBJ_BG, OBJ_ONLY, PB_T50_RS) are shown in Table 1. For each variant, we conducted experiments 8 times with different random seeds and reported the best and the average scores, and it is observed that: (i)Holistically, with only 1.05M, 4.7% parameters of Full fine-tuning paradigm, the performance of our block is improved by +0.7% and +11.3% compared with that of standard Transformer under Full tuning protocol on ReCon and Point-MAE, respectively. (ii) Compared with the current best methods IDPT and DAPT, our PPT also gets better results than both of them. Specifically, we achieve 2.4% and 4.5% performance improvement compared with DAPT and IDPT respectively. That is, with PPT on both ReCon and Point-MAE achieve state-of-the-art performance.
3D Synthetic Object Recognition
ModelNet (Wu et al. 2015) is one of the most classical datasets for synthetic 3D object recognition. It contains 12K meshed 3D CAD objects of 40 (ModelNet40) or 10 (ModelNet10) categories. We conducted the evaluation on the ModelNet40 dataset, including fine-tuning and few-shot learning. During training and testing, we use Scale&Translate as data augmentation in training following (Qi et al. 2017a, b). The results are shown in Table 1 and Table 2, respectively. We report two different settings with 1k points and 8k points respectively. It can observed that (i)With only a few trainable parameters, compared with Full tuning protocol, ReCon gains 2.1% performance improvement with our PPT. Point-MAE also gets results on par with the Full tuning with less than 5% of the parameters trainable. (ii)Compared with the current best methods IDPT and DAPT, with our PPT, we achieve 1.26% and 0.32% improvements higher than these two respectively. Point-MAE with PPT achieves the best result on ModelNet40 with 93.88% accuracy. And ReCon with PPT gains +1.3% and +0.8% performance improvements, showing the effectiveness of our positional prompt tuning method.
@@
| PointNet (Qi et al. 2017a) | CVPR 17 | - | 80.39 | 83.7 |
| PointNet++ (Qi et al. 2017b) | NeurIPS 17 | - | 81.85 | 85.1 |
| DGCNN (Wang et al. 2019) | TOG 19 | - | 82.33 | 85.2 |
@@
| OcCo (Wang et al. 2021) | ICCV 21 | 27.09 | 83.42 | 85.1 |
| Point-BERT (Yu et al. 2022) | CVPR 22 | 27.09 | 84.11 | 85.6 |
| Point-MAE (Pang et al. 2022) | ECCV 22 | 27.06 | 84.19 | 86.1 |
| ACT(Dong et al. 2023) | ICLR 23 | 27.06 | 84.66 | 86.1 |
| ReCon(Qi et al. 2023a) | ICML 23 | 27.06 | 84.66 | 86.4 |
| with Parameter-Efficient Supervised Finetuning | ||||
| Point-MAE (Pang et al. 2022) (baseline) | ECCV 22 | 27.06 | 84.19 | 86.1 |
| + IDPT (Zha et al. 2023) | ICCV 23 | 5.69 | 83.79 | 85.7 |
| + DAPT (Zhou et al. 2024) | CVPR 24 | 5.65 | 84.01 | 85.7 |
@@
| ReCon (Qi et al. 2023a) (baseline) | ICML 23 | 27.06 | 84.52 | 86.1 |
| + IDPT (Zha et al. 2023) | ICCV 23 | 5.69 | 83.66 | 85.7 |
| + DAPT (Zhou et al. 2024) | CVPR 24 | 5.65 | 83.87 | 85.7 |
| + PPT (ours) | - | 5.62 | 84.23 | 85.6 |
Few-shot Learning
We conducted experiments of few-shot learning on the ModelNet40 (Wu et al. 2015) dataset. We considered settings of n-way k-shot manner following previous works (Pang et al. 2022; Qi et al. 2023a), where n 5, 10 and k 10, 20. As shown in Table 2, it can be observed that our PPT can achieve performance enhancement in most cases compared with IDPT and DAPT, which also indicates the effectiveness of our method.
3D Part Segmentation
To evaluate the geometric understanding performance within objects, we conducted the part segmentation experiment on ShapeNetPart (Yi et al. 2016). Specifically, we concatenate the cross-modal feature into the global feature and use the same segmentation head as Point-MAE for a fair comparison. From Table 3, it can be observed that our method gets results on par with from scratch baseline on both Cls. mIoU and Inst. mIoU. Besides, PPT also outperforms the PEFT counterparts IDPT (Zha et al. 2023) and DAPT (Zhou et al. 2024) 0.06% and 0.46% Cls. mIoU, respectively. It shows that the cross-modal global knowledge from cross-modal pretraining can still play a certain role in part segmentation.
| Methods | Params. (M) | OBJ_BG | OBJ_ONLY | PB_T50_RS |
@@
Discussions
Position is all your need in 3D Transformer?
@@
Figure 4 illustrates this very simple architecture, and its classification performance on ScanObjectNN is shown in Table 4. With only 0.8M parameters, PosNet achieved remarkable performance, further validating our theoretical assumption that the global understanding provided by positional encoding is crucial in 3D Transformers.
Conclusions
In this study, we rethink the role of positional encoding in 3D representation learning and fine-tuning. We argue that using positional encoding in point Transformer-based methods serves to aggregate multi-scale features of point clouds. Additionally, we explore parameter-efficient fine-tuning (PEFT) through the lens of prompts and adapters, introducing a straightforward yet effective method called PPT for point cloud analysis. PPT incorporates increased patch tokens and trainable positional encoding while keeping most pre-trained model parameters frozen. Extensive experiments validate that PPT is both effective and efficient.
Appendix A Additional Experiments
Implement Details
To ensure a fair comparison, we employed identical experimental settings to the default fine-tuning method for each baseline. This entails freezing the weights of the pre-trained point cloud backbone and solely updating the newly inserted parameters during training. All experiments are conducted on a single GeForce RTX 3090.
Ablation Studys
We have done several ablation experiments on our method. In Table 7 we replace the positional embedding layer in the PPT with several classic forms. The means position encoding is frozen during the tuning process. As we can see in the table, the results show that such MLP structures as position embedding layers, as we mentioned before, perform better than others, which also proves that the point Transformer with MLP embedding layers serves as a multi-scale feature extractor. And it’s surprising that tuning with only the first layer receives MLP-embedded positional information can achieve relatively high performance, even better than all other forms. We also set an ablation study on the number of patch token groups K. Specifically, we set the number K range from 1 to 3, and K=2 turns out to be the best choice, demonstrating the efficiency of the extra prompt tokens, and also, too many prompt token groups will interference the process of fine-tuning. Nevertheless, the result of K=3 which is relatively higher than the result of K=1 also shows the effectiveness of the extra token prompts edgewise.

### v2

@@ 3.2. Better way of prompt and adapter tuning?
With point cloud as the only modality of input, it is difficult for 3D parameter-efficient fine-tuning tasks to get prompts that are rich in linguistic significance, like a well-designed prefix prompt in natural language, which can remarkably boost the performance of generative tasks in text modality (Lester et al., 2021). VPT (Jia et al., 2022) introduces a set of trainable embeddings into the input of the Transformer to solve this problem in the vision domain, and both DAPT (Zhou et al., 2024) and IDPT (Zha et al., 2023) adopt a similar manner. Simultaneously considering Adapter and Prompt, they both generate extra prompt tokens from an adapter-structure network and add these prompt tokens into the Transformer input.
By doing so, they acquire dynamic prompts rather than static ones in VPT. As the static prompts always need a process of manual selection, the generated dynamic prompts are a compromise of this process. Nevertheless, these kinds of prompts are actually an abstraction of the existing feature embedding. Thus the input of Transformer layers with extra prompt inserted and the Transformer output can be formulated as:
@@
Where means class token, represents the extra prompts abstracted from patch embedding . attn and F represent the FFN and attention layer inside the Transformer blocks. The and represent trainable and frozen parameters, respectively. Thus, the information of prompts is highly homogeneous, which is an aggregation of existing tokens, rather than the newly added language prompts in the NLP domain. We then think about whether there are some other forms of prompts that can both be dynamically adjusted and, in the meantime, be of physical significance. And it comes to our mind that the patch tokens themselves, which are down-sampled and clustered, can be used as extra tokens for prompting. Different from prompts that are yielded from patch token embedding, we unfreeze a part of the patch encoder to introduce variety into patch token embedding. In that case, with this form of prompts, the input of Transformer layers can be formulated as:
@@
represents the added trainable adapter for more thorough feature aggregation and adjustment. In this way, both the extra sampled patch token prompts and the original patch token embedding are altered for information integration. In this manner of prompts and adapters, combined with the aforementioned positional encoding, we carried out extensive experiments, and our method achieved excellent results in all of these experiments, which will be discussed in subsequent chapters.
@@
| DGCNN (Wang et al., 2019) | TOG 19 | 31.6 2.8 | 40.8 4.6 | 19.9 2.1 | 16.9 1.5 |
| OcCo (Wang et al., 2021) | ICCV 21 | 90.6 2.8 | 92.5 1.9 | 82.9 1.3 | 86.5 2.2 |
@@
| OcCo (Wang et al., 2021) | ICCV 21 | 94.0 3.6 | 95.9 2.3 | 89.4 5.1 | 92.4 4.6 |
| Point-BERT (Yu et al., 2022) | CVPR 22 | 94.6 3.1 | 96.3 2.7 | 91.0 5.4 | 92.7 5.1 |
| Point-MAE (Pang et al., 2022) | ECCV 22 | 96.3 2.5 | 97.8 1.8 | 92.6 4.1 | 95.0 3.0 |
| Point-M2AE (Zhang et al., 2022) | NeurIPS 22 | 96.8 1.8 | 98.3 1.4 | 92.3 4.5 | 95.0 3.0 |
| ACT (Dong et al., 2023) | ICLR 23 | 96.8 2.3 | 98.0 1.4 | 93.3 4.0 | 95.6 2.8 |
| VPP (Qi et al., 2023b) | NeurIPS 23 | 96.9 1.9 | 98.3 1.5 | 93.0 4.0 | 95.4 3.1 |
| I2P-MAE (Zhang et al., 2023d) | CVPR 23 | 97.0 1.8 | 98.3 1.3 | 92.6 5.0 | 95.5 3.0 |
| ReCon (Qi et al., 2023a)(baseline) | ICML 23 | 97.3 1.9 | 98.9 1.2 | 93.3 3.9 | 95.8 3.0 |
| with Parameter-Efficient Supervised Fine-tuning | |||||
| ReCon w/ IDPT∗ (Zha et al., 2023) | ICCV 23 | 96.92.4 | 98.30.7 | 92.84.0 | 95.53.2 |
| ReCon w/ DAPT∗ (Zhou et al., 2024) | CVPR 24 | 95.62.8 | 97.71.6 | 91.94.1 | 94.63.5 |
| ReCon w/ PPT (Ours) | ACMMM 25 | 97.0 2.7 | 98.71.6 | 92.25.0 | 95.62.9 |
3.3. Pipeline of PPT
@@
The pipeline structure of our PPT is shown in fig.˜3(a). The patch encoder of the pre-trained model is already empowered to abstract a high concentration of semantic information from the input point cloud, so we unfreeze part of the patch encoder to provide variety for patch tokens and prompt tokens. Briefly, we resample the input point clouds as prompts and concatenate them with the input patch tokens before sending them into the Transformer, and trainable positional embedding layers are concatenated together as the positional information for both the prompts and the patch tokens. Before inputting it into each Transformer layer, a patch adapter layer and a prompt adapter layer are adopted as a dynamic adjustment module. There are also adapter layers inside every Transformer block, specifically between the attention layer and the FFN. As the pre-trained parts of the Transformer layers are untrainable during the tuning process, an adapter inside the block effectively integrates information and dynamically adjusts it to improve performance.
3.4. Algorithm Pseudo Coding
In algorithm˜1, we report the pseudo-code of our algorithm. As we can see from the fairly simple code, the overall implementation process is concise but very efficient.
4. Experiments
4.1. 3D Real-World Object Recognition
@@ 4. Experiments
ScanObjectNN (Uy et al., 2019) is one of the most challenging 3D datasets, which covers 15K real-world objects from 15 categories. For a fair comparison, we report the results with and without the voting strategy (Liu et al., 2019) separately. The results of the object classification tasks of ScanObjectNN (OBJ_BG, OBJ_ONLY, PB_T50_RS) are shown in table˜1. For each variant, we conducted experiments 8 times with different random seeds and reported the best and the average scores, and it can be observed that: (i)Holistically, with only 1.05M, 4.7% parameters of Full fine-tuning paradigm, the performance of our block is improved by +0.7% and +11.3% compared with that of standard Transformer under Full tuning protocol on ReCon and Point-MAE, respectively. (ii) Compared with the current best methods, IDPT and DAPT, our PPT also gets better results than both of them. Specifically, we achieve 2.4% and 4.5% performance improvement compared with DAPT and IDPT, respectively. That is, with PPT on both ReCon and Point-MAE achieve state-of-the-art performance.
@@
| PointNet (Qi et al., 2017a) | CVPR 17 | - | 80.39 | 83.7 |
| PointNet++ (Qi et al., 2017b) | NeurIPS 17 | - | 81.85 | 85.1 |
| DGCNN (Wang et al., 2019) | TOG 19 | - | 82.33 | 85.2 |
@@
| OcCo (Wang et al., 2021) | ICCV 21 | 27.09 | 83.42 | 85.1 |
| Point-BERT (Yu et al., 2022) | CVPR 22 | 27.09 | 84.11 | 85.6 |
| Point-MAE (Pang et al., 2022) | ECCV 22 | 27.06 | 84.19 | 86.1 |
| ACT(Dong et al., 2023) | ICLR 23 | 27.06 | 84.66 | 86.1 |
| ReCon(Qi et al., 2023a) | ICML 23 | 27.06 | 84.66 | 86.4 |
| with Parameter-Efficient Supervised Fine-tuning | ||||
| Point-MAE (Pang et al., 2022) (baseline) | ECCV 22 | 27.06 | 84.19 | 86.1 |
| + IDPT (Zha et al., 2023) | ICCV 23 | 5.69 | 83.79 | 85.7 |
| + DAPT (Zhou et al., 2024) | CVPR 24 | 5.65 | 84.01 | 85.7 |
@@
| ReCon (Qi et al., 2023a) (baseline) | ICML 23 | 27.06 | 84.52 | 86.1 |
| + IDPT (Zha et al., 2023) | ICCV 23 | 5.69 | 83.66 | 85.7 |
| + DAPT (Zhou et al., 2024) | CVPR 24 | 5.65 | 83.87 | 85.7 |
| + PPT (ours) | ACMMM 25 | 5.62 | 84.23 | 85.6 |
4.2. 3D Synthetic Object Recognition
ModelNet (Wu et al., 2015) is one of the most classical datasets for synthetic 3D object recognition. It contains 12K meshed 3D CAD objects of 40 (ModelNet40) or 10 (ModelNet10) categories. We conducted the evaluation on the ModelNet40 dataset, including fine-tuning and few-shot learning. During training and testing, we use Scale&Translate as data augmentation in training following (Qi et al., 2017a, b). The results are shown in table˜1 and table˜2, respectively. We report two different settings with 1k points and 8k points, respectively. It can be observed that (i) With only a few trainable parameters, compared with Full tuning protocol, ReCon gains 2.1% performance improvement with our PPT. Point-MAE also gets results on par with the Full tuning with less than 5% of the parameters trainable. (ii) Compared with the current best methods, IDPT and DAPT, with our PPT, we achieve 1.26% and 0.32% improvements higher than these two, respectively. Point-MAE with PPT achieves the best result on ModelNet40 with 93.88% accuracy. And ReCon with PPT gains +1.3% and +0.8% performance improvements, showing the effectiveness of our positional prompt-tuning method.
| Num K | #TP (M) | PB_T50_RS |
| 1 | 0.9M | 87.40 |
| 2 | 1.1M | 89.52 |
| 3 | 1.4M | 88.06 |
4.3. Few-shot Learning
We also conduct experiments of few-shot learning on the ModelNet40 (Wu et al., 2015) dataset. We consider settings of n-way k-shot manner following previous works (Pang et al., 2022; Qi et al., 2023a), where n 5, 10 and k 10, 20. As shown in table˜2, it can be observed that our PPT can achieve performance enhancement in most cases compared with IDPT (Zha et al., 2023) and DAPT (Zhou et al., 2024), which also indicates the effectiveness of our method.
4.4. 3D Part Segmentation
To evaluate the geometric understanding performance within objects, we conducted the part segmentation experiment on ShapeNetPart (Yi et al., 2016). Specifically, we concatenate the cross-modal feature into the global feature and use the same segmentation head as Point-MAE for a fair comparison. From table˜3, it can be observed that our method gets results on par with the from scratch baseline on both Cls. mIoU and Inst. mIoU. Besides, PPT also outperforms the PEFT counterparts IDPT (Zha et al., 2023) and DAPT (Zhou et al., 2024) by 0.06% and 0.46% Cls. mIoU, respectively. It shows that the cross-modal global knowledge from cross-modal pretraining can still play a certain role in part segmentation.
4.5. Ablation Studys
We conduct ablation experiments to demonstrate the effectiveness of each module in our method. Details and results are provided as follows.
Ablation study on position embedding forms. In table˜5, we replace the positional embedding layer in the PPT with several classic forms. As we can see in the table, the results show that such MLP structures as position embedding layers, as we mentioned before, perform better than others, which also proves that the point Transformer with MLP embedding layers serves as a multi-scale feature extractor. It’s surprising that tuning with only the first layer receives MLP-embedded positional information can achieve relatively high performance, even better than all other forms.
Ablation on group number K. We conduct an ablation study on the number of patch token groups K, ranging from 1 to 3. The best choice is K=2, showing the effectiveness of extra prompt tokens. Too many prompt token groups can interfere with fine-tuning. However, K=3, though higher than K=1, still demonstrated the benefit of extra token prompts.
| Methods | Params. (M) | OBJ_BG | OBJ_ON | PB_T50_RS |
@@ 4.5. Ablation Studys
5. Discussions
5.1. Position is all your need in 3D Transformer?
@@
fig.˜4 illustrates this very simple architecture, and its classification performance on ScanObjectNN is shown in table˜6. With only 0.8M parameters, PosNet achieved remarkable performance, further validating our theoretical assumption that the global understanding provided by positional encoding is crucial in 3D Transformers.
5.2. Attention Visualization in Position Encodings
We also report the attention score visualization of either position embedding or patch tokens as the input, respectively. Taking the 3D coordinates of patch centers as the input of position embedding, the attention scores of the position embedding input are more holistic, as shown in the top row of fig.˜5. That means the position embedding features are concerned with more global information. Meanwhile, the patch token embedding, with all neighbor points around the cluster centers as input, gathers together their attention gathered together at some points, as the local counterpart of position embedding inputs. Concerning both global and local features, the point cloud Transformer architectures deal with complex tasks with a simple MLP as the position encoding module and even get better results than adding those well-designed position embedding methods together, as reported in table˜5.
6. Limitations & Future Work
Although we want to underscore the significance of positional encoding within 3D Transformer and have experimentally validated the use of positional encoding as a form of prompt tuning in classic 3D self-supervised pre-training works, such as Point-MAE (Pang et al., 2022) and ReCon (Qi et al., 2023a), we have yet to conduct experimental validation on larger datasets, such as Objaverse (Deitke et al., 2023). Our primary future work will be to experiment with positional prompt tuning on a wider range of 3D foundational models, such as OpenShape (Liu et al., 2023) and ReCon++ (Qi et al., 2024), and to expand its application to a broader spectrum of tasks, including zero-shot learning (Xue et al., 2023; Liu et al., 2023; Qi et al., 2023a, 2024; Xie et al., 2024; Lin et al., 2025), 3D AIGC (Nichol et al., 2022; Qi et al., 2023b; Ma et al., 2024a), 3D large language models (Qi et al., 2024; Dong et al., 2024; Tang et al., 2025; Xiong et al., 2025), and robotics (Qi et al., 2025; Zhang et al., 2025; Jia et al., 2025; He et al., 2025), among others (Dong et al., 2022; Peng et al., 2024; Zhang et al., 2023b; Yu et al., 2024). There are also plenty of new methods related to 3D representation learning or parameter-efficient fine-tuning (Wang et al., 2024b; Kang et al., 2024; Liu et al., 2025; Sun et al., 2025, 2024a), which we will conduct experiments on.
7. Conclusions
In this paper, we revisit the role of position encoding in 3D Transformer models, highlighting its simplicity and importance as a high-dimensional element. By combining position encoding with the patch encoder, we capture multi-scale information. This, alongside the sequential Transformer, creates a comprehensive framework that integrates both local features from patches and global features from center points through position encoding. With just a few parameters, the position embedding module aligns well with PEFT tasks, making it ideal for fine-tuning. Extensive experiments validate that PPT is both effective and efficient.
Acknowledgments
This work was sponsored by the National Natural Science Foundation of China No. 62472349, the Fundamental Research Funds for the Central Universities No. xxj032023020, and CAAI-CANN Open Fund, developed on OpenI Community.
Appendix A Implement Details
A.1. Training recipes